### v1

@@ II.1 Open questions
The majority of solar neutrino fluxes have now been measured, in some cases with a precision equivalent to or better than theoretical predictions. Improved precision in several cases can offer further insight. A more precise measurement of the flux of neutrinos from the CNO cycle (Fig. 2) would test important aspect of the SSM including, in particular, the initial abundance of heavy elements [2007A&A...463..755C, 2011ApJ...743...24S, 2022A&A...667L...2K] and the redistribution of the CNO elements from out-of-equilibrium burning at the onset of the main sequence. Solar neutrino flux measurements can constrain physical processes in stars such as chemical mixing. The fluxes are sensitive to solar/stellar model inputs, including the radiative opacities (see Sect. LABEL:sec:Opacities). Measurements of the 8B and 7Be fluxes can offer a handle on the temperature of the solar core, and other environmental factors. A percent-level measurement of the pp flux would provide a test of the luminosity constraint, further probing solar power generation mechanisms. The one branch of the pp-chain that yet remains undetected, the hep flux, has both the highest energy but also the lowest flux of neutrinos from the Sun. The spectrum extends beyond that of the 8B neutrinos, offering a small window for unequivocal observation. Limits have been placed on this flux [2020PhRvD.102f2006A], but a definitive observation will likely require a large, next-generation detector such as Hyper-Kamiokande (HyperK) [2018arXiv180504163H] or the Deep Underground Neutrino Experiment (DUNE) [2019PhRvL.123m1803C]. Finally, the observation of monochromatic neutrinos produced by electron capture reactions on 13N, 15O and 17F (which we refer to as ecCNO neutrinos) [1990PhRvD..41.2964B, 2004PhRvC..69a5801S, 2015PhLB..742..279V] would provide important new physics. They can be used as probes of the metal content of the solar core. They provide a measure of the electron neutrino survival probability at specific neutrino energies MeV within the broad transition region between vacuum and matter-enhanced oscillations, where currently we have no constraints. This would constitute a new test of the large-mixing-angle (LMA) MSW flavor oscillation paradigm. The detection of this subdominant component of the CNO-cycle is extremely difficult but it could be within reach of future very large ultra-pure liquid scintillator detectors [2015PhLB..742..279V].
@@
There is also opportunity to leverage solar neutrinos to understand the interaction of neutrinos with matter. The effect of MSW oscillations has a significant impact on the observed solar neutrino spectrum: with vacuum oscillation dominating at low energy, below approximately 1 MeV, where the survival probability is roughly one half, while matter effects further suppress the flux to a survival probability of roughly one third above approximately 5 MeV. The transition region between these two regimes offers an extremely sensitive probe of the details of the interactions of neutrinos with matter, including the potential to search for new physics such as sterile neutrinos, or non-standard interactions, by looking for distortions to the expected spectral shape.
@@
The same theory of matter effects predicts a small regeneration of electron neutrinos during the night time, as they propagate through the bulk of the earth. This so-called “day/night effect” has been sought after by both Super-Kamiokande (SuperK) and the Sudbury Neutrino Observatory (SNO) [2014PhRvL.112i1805R, 2002PhRvL..89a1302A], but a significant observation is still limited by statistics. Future data from HyperK or DUNE may be needed to confirm our understanding of this prediction of the MSW effect.
Mild tension lingers between measurements of the mass splitting parameters, between solar neutrino experiments and terrestrial data sensitive to the same parameter, from the KamLAND reactor experiment [Esteban2020TheFO]. Further study of solar neutrinos could allow resolution of this question, potentially offering new insights.
@@ II.1.1 The solar composition problem
Solar photospheric abundances, determined with spectroscopic techniques, are a fundamental input for the construction of SSMs, but also they are used as input to nearly all models in astrophysics, including stellar evolution, protoplanetary disks, and galactic chemical evolution. The development of three dimensional radiation hydrodynamic models of the solar atmosphere, of techniques to study line formation under non-local thermodynamic conditions and the improvement in atomic properties (e.g. transition strengths) have led since 2001 to a complete revision of solar abundances. Initial results based on these more sophisticated methods pointed towards a markedly lower solar metallicity [2009ARA&A..47..481A], in particular regarding CNO elements, than results from the 90s based on older techniques [1993oee..conf...15G, 1998SSRv...85..161G].
@@
The degeneracy between solar composition and opacities can be broken using CNO solar neutrinos, e.g., following the methodology developed in 2008ApJ...687..678H. Such a study was recently carried out by the Borexino collaboration [2020Natur.587..577B, 2022PhRvL.129y2701A, 2023PhRvD.108j2005B]. The Borexino measurement of the CN-cycle neutrino flux (the 13N and 15O fluxes) were used to determine the C+N core abundance. Results show a tension with LZ metallicity determinations, while being in better agreement with HZ mixtures. While the error budget is presently dominated by the uncertainty of the Borexino CNO neutrino measurement, a significant contributor to the error () is nuclear, due to uncertainties in , , and . The interpretation of future improved CNO neutrino measurements will be impacted, unless these nuclear physics uncertainties are reduced.
@@ II.1.2 The Gallium anomaly
The precisely known rate of electron capture on 71Ge to 71Ga, for which the half-life is 11.43(3) d [1985PhRvC..31..666H], places an important constraint on the neutrino absorption rate [2023PhLB..84237983G, PhysRevC.108.035502]. The allowed matrix element that governs neutrino capture to the ground state of 71Ge can be extracted to a precision of 1% from the electron capture rate. The ground-state transition by itself generates a significant anomaly: inclusion of the excited-state contributions doubles the effect. In Appendix A we describe some of the details of both the 1997 extraction of the allowed (Gamow Teller) strength by 1997PhRvC..56.3391B and the recent update of PhysRevC.108.035502, ELLIOTT2024104082.
@@ II.1.3 The Boron-8 neutrino spectrum
In SF II it was decided to recommend the spectrum calculated by Winter et al. [2003PhRvL..91y2501W, 2006PhRvC..73b5503W, Win:2007] based on their measurement of the alpha spectrum from . This spectrum showed excellent agreement with an independent experiment [2006PhRvC..73e5802B]. Both radiative and recoil-order corrections were included by Winter et al. The radiative corrections are relatively small, due to a cancellation between the real and virtual contributions [1995PhRvD..52.5362B]. The recoil-order corrections are dominated by the weak magnetism part, which has been deduced from measurements of the analog electromagnetic decays [1995PhRvC..51.2778D].
Since then three additional measurements [2011PhRvC..83f5802K, 2012PhRvL.108p2502R, 2023PhRvC.107c2801L] of the -decay alpha spectrum have been performed. All three find that the peak of the alpha spectrum appears about 20 keV lower than determined by Winter et al. Overall, these new measurements yield differences in the neutrino spectrum of 5% below MeV, where the contribution becomes small compared to the hep component. Figure 4 shows the 8B spectral uncertainties on the scale of the hep spectrum. With resolution effects considered, it will be difficult to extract the hep flux from solar neutrino measurements without further reduction in 8B uncertainties.
@@ II.2 Experimental program on solar neutrinos
A broad range of technology can be used to interrogate solar neutrinos. Radiochemical experiments utilizing the interaction on either chlorine or gallium were the first to identify the mystery that became known as the Solar Neutrino Problem. Large monolithic detectors have achieved great success in real-time measurements – from the water Cherenkov detectors such as Kamiokande, SNO, and SuperK, which could use the unique topology of Cherenkov light to point back to the neutrinos’ origin, to liquid scintillator detectors such as KamLAND and Borexino, whose high light yield and low threshold allow for precision spectroscopy.
@@ II.2.1 Super-Kamiokande and Hyper-Kamiokande
The first real-time detection of solar neutrinos was achieved by the Kamiokande experiment [1989PhRvL..63...16H], which detected the neutrinos via elastic scattering (ES) of electrons, a process in which the outgoing electron’s direction is highly correlated to that of the incoming neutrino. This allowed Kamiokande to directly point back to the neutrinos’ origin: our Sun. The successor experiment, Super-Kamiokande, is a 50-kton volume of pure water, surrounded by over 10,000 photon detectors, which has achieved an unparalleled program of neutrino observations and other physics over its several decades of operation. This program has included the highest-precision measurement of the ES signal from solar neutrino interactions, as well as sensitive searches for the day/night distortions of the spectral shape [2016PhRvD..94e2010A]. Now filled with a gadolinium additive to enhance neutron capture, SuperK’s primary current focus is the search for the Diffuse Supernova Neutrino Background (DSNB) via inverse beta decay [2023ApJ...951L..27H]. It also continues its atmospheric neutrino program: SuperK measurements made 25 years ago demonstrated the neutrino oscillations were responsible for the puzzling zenith-angle-dependence of this flux [PhysRevLett.81.1562].
@@ II.2.2 Sudbury Neutrino Observatory and SNO+
SNO’s unique use of heavy water as a target medium offered the ability to detect solar neutrinos via two additional interactions besides ES. In the charged current (CC) interaction an electron neutrino interacts with the neutron in the deuteron, producing an electron and a proton. This interaction path is sensitive only to electron-flavor neutrinos at the few-MeV energy scale of solar neutrinos, providing a measurement of the pure flux. SNO also had access to the neutral current (NC) interaction, in which a neutrino of any flavor interacts with the deuteron, breaking it apart into its constituent nucleons. Being equally sensitive to all active flavors, the NC interaction yields a measurement of the total, flavor-blind neutrino flux. It was this capability that allowed SNO to resolve the Solar Neutrino Problem, and demonstrate that the produced in the Sun were changing flavor prior to detection [2001PhRvL..87g1301A, 2002PhRvL..89a1301A, 2004PhRvL..92r1301A].
@@
SNO ceased data taking in 2006. Analysis of the data set has continued, and produced a number of new results, including constraints on non-standard effects such as Lorentz violation and neutrino decay [2018PhRvD..98k2013A, 2019PhRvD..99c2013A], searches for hep neutrinos and the diffuse supernova neutrino background (DSNB) [2020PhRvD.102f2006A], and studies of neutron production from cosmogenic muons and from atmospheric neutrino interactions [2019PhRvD..99k2007A, 2019PhRvD.100k2005A].
@@
After decommissioning, the detector was repurposed for the SNO+ experiment, in which the target material was replaced with a pure organic liquid scintillator (LS): linear alkyl benzene (LAB), loaded with 2.2 g/L of the secondary fluor, PPO. The high light yield of this scintillator, coupled with the location in SNOLAB –at 6-km water equivalent, one of the deepest underground laboratories in the world– and an extremely well-understood detector, allow for a range of high-precision measurement programs. In a preliminary water phase, SNO+ demonstrated detection of neutron captures on hydrogen [PhysRevC.102.014002], an impressive technical achievement in an unloaded water detector; a low background measurement of 8B solar neutrinos [PhysRevD.99.012012]; several searches for invisible modes of nucleon decay [PhysRevD.99.032008, PhysRevD.105.112012]; and detection of antineutrinos from reactors over 200 km away [PhysRevLett.130.091801]. These results demonstrate both the low backgrounds and the technical capabilities of the detector, which will facilitate the future program. In a “partial-fill” stage, in which the upper half of the detector was filled with LS while the lower half still contained water, the first demonstration of event-level direction reconstruction was achieved for 8B neutrinos in a scintillator detector [PhysRevD.109.072002]. This was facilitated by a lower loading of PPO at that time, which results in lower scintillation yield and a slower time profile, effectively enhancing the Cherenkov component, which can be leveraged for directional information.
@@
Now fully filled with LS, SNO+ is the deepest, largest operating LS detector in the world. The future program will include measurements of several solar neutrino fluxes, as well as the 8B neutrino energy spectrum.
@@ II.2.3 Borexino
Water Cherenkov experiments such as Kamiokande, SuperK and SNO provided the first real-time measurement of solar neutrinos but their relatively high energy threshold made them sensitive only to a small fraction of the solar neutrino flux. To study in real time the bulk of solar neutrino emission a different detector technology was required. The Borexino experiment, located deep underground at the INFN Laboratori Nazionali del Gran Sasso, used an organic LS target made of pseudocumene with PPO to detect the ES of solar neutrinos off electrons. The high light yield of the LS made it possible to significantly lower the energy threshold, but given the small signal rate and the lack of a clear signature to separate it from the background (such as the direction indicated by the emission of Cherenkov radiation), the experiment required a long preparatory phase to develop the most advanced techniques to suppress the background, especially the one due to radioactive contamination of the LS itself.
@@
The detector’s extreme radiopurity was key to its success, and over its 14 years of data taking, ending in October 2021, Borexino proved itself capable of covering the entire solar neutrino spectrum. At the beginning of its data taking in 2007 the level of 238U and 232Th contamination in Borexino were lower than and respectively [2008PhLB..658..101B], paving the way for the first measurement of the 7Be sub-MeV solar neutrinos, followed by a low-threshold measurement of the 8B flux and a first indication of the pep neutrinos. After a purification campaign that further reduced the LS contamination, Borexino aimed at improving the accuracy of its results in the measurement of the pp-chain solar neutrinos. The flux of neutrinos produced in the pp fusion process was first measured in [2014Natur.512..383B] and later improved along with all the other fluxes produced throughout the pp-chain, with the only exception of hep neutrinos [2018Natur.562..505B, 2019PhRvD.100h2004A, 2020PhRvD.101f2001A]: the pp flux was determined to an uncertainty of 11%, improving the neutrino-based estimate of solar luminosity, while 7Be neutrinos were measured with an uncertainty of , half that of the SSM prediction. Furthermore, applying a constraint on the flux of CNO neutrinos derived from the SSM predictions, the signal of pep neutrinos was established with a significance larger than for the first time, and the 8B flux was measured with a threshold as low as 3 MeV.
@@
After completing the investigation of the pp-chain neutrinos, Borexino reported the first detection of solar neutrinos produced in the CNO-cycle [2020Natur.587..577B], providing the first direct evidence of the occurrence of that process in stars. Profiting from a larger exposure and a better understanding of the radioactive backgrounds enabled by the unprecedented thermal stability of the detector, Borexino further improved its measurement of the CNO neutrino flux [2022PhRvL.129y2701A, 2023PhRvD.108j2005B], where the new CNO result was used in combination with existing solar neutrino results to probe solar composition.

### v2

@@ II.1 Open questions
The solar neutrino spectrum has been measured for energies above 233 keV, with some fluxes determined to a precision equivalent to or better than theoretical predictions. However, improved measurements can offer further insight on several aspects. For instance, more precise determinations of the fluxes of neutrinos from the CNO cycle (Fig. 2) could determine the contemporary abundances of N and O in the solar core [2007A&A...463..755C, 2011ApJ...743...24S, 2022A&A...667L...2K]. While these abundances differ significantly from the primordial values that are input into the SSM –the primordial abundances are altered by out-of-equilibrium CNO-cycle burning that occurs in the early Sun [Roxburgh85]– this burning redistributes the C, N, and O, but does not alter the net number of such nuclei. Consequently, CNO neutrino flux measurements might allow one to experimentally address the solar composition problem discussed below.
Solar neutrino flux measurements can also constrain physical processes in stars such as chemical mixing. The fluxes are sensitive to solar/stellar model inputs, including the radiative opacities (see Sect. LABEL:sec:Opacities). Measurements of the neutrino fluxes from 8B and 7Be can offer a handle on the temperature of the solar core, and other environmental factors [1996PhRvD..53.4202B].
A measurement of the pp flux with percent-level precision would provide a test of the luminosity constraint, further probing solar power generation mechanisms [2002PhRvC..65b5801B, 2021JPhG...48a5201V]. The one branch of the pp chain that yet remains undetected, the hep flux, has both the highest energy but also the lowest flux of neutrinos from the Sun. The spectrum extends beyond that of the 8B neutrinos, offering a small window for unequivocal observation. Limits have been placed on this flux [2020PhRvD.102f2006A], but a definitive observation will likely require a large, next-generation detector such as Hyper-Kamiokande (HyperK) [2018arXiv180504163H] or the Deep Underground Neutrino Experiment (DUNE) [2019PhRvL.123m1803C].
Finally, the observation of monochromatic neutrinos produced by electron capture reactions on 13N, 15O and 17F (which we refer to as ecCNO neutrinos) [1990PhRvD..41.2964B, 2004PhRvC..69a5801S, 2015PhLB..742..279V] could be rewarding in terms of physical implications. Indeed, they can be used as probes of the metallicity of the solar core. Moreover, they provide a measure of the electron neutrino survival probability at specific neutrino energies MeV within the broad transition region between vacuum and matter-enhanced oscillations, where currently we have no constraints. This would constitute a new test of the large-mixing-angle (LMA) MSW flavor oscillation paradigm. The detection of this subdominant component of the CNO-cycle is extremely difficult but it could be within reach of future very large ultra-pure liquid scintillator detectors [2015PhLB..742..279V].
@@
Solar neutrinos provide an opportunity to understand the interaction of neutrinos with matter. The effect of MSW oscillations has a significant impact on the observed solar neutrino spectrum: with vacuum oscillation dominating at low energy, below approximately 1 MeV, where the survival probability is roughly one half, while matter effects further suppress the flux to a survival probability of roughly one third above approximately 5 MeV. The transition region between these two regimes offers an extremely sensitive probe of the details of the interactions of neutrinos with matter, including the potential to search for new physics such as sterile neutrinos, or non-standard interactions, by looking for distortions to the expected spectral shape [2019arXiv190700991B, 2020JHEP...02..038B, 2023JHEP...08..032C].
@@
The same theory of matter effects predicts a small regeneration of electron neutrinos during the night time as they propagate through the bulk of the earth. This so-called “day/night effect” has been sought after by both Super-Kamiokande (SuperK) and the Sudbury Neutrino Observatory (SNO) [2014PhRvL.112i1805R, 2002PhRvL..89a1302A], but a significant observation is still limited by statistics. Future data from HyperKamiokande or DUNE may be needed to confirm our understanding of this prediction of the MSW effect.
The slight tension between measurements of the mass splitting parameters, in solar neutrino experiments and terrestrial data sensitive to the same parameter, from the KamLAND reactor experiment has now disappeared with the addition of new data on day-night effects and spectral shapes [Esteban2020TheFO, 2018PrPNP.102...48C]. Further tightening of the oscillation parameters can be expected from JUNO, which will obtain highly precise data on both the solar mass splitting and the mixing angle [2016JPhG...43c0401A].
@@ II.1.1 The solar composition problem
Solar photospheric abundances, determined with spectroscopic techniques, are a fundamental input for the construction of SSMs, but also they are used as input to nearly all models in astrophysics, including stellar evolution, protoplanetary disks, and galactic chemical evolution. Over the last two decades, the development of three dimensional radiation hydrodynamic models of the solar atmosphere [2009MmSAI..80..711L, 2000A&A...359..729A] and of techniques to study line formation under non-local thermodynamic conditions [2005ARA&A..43..481A], together with improved descriptions of atomic properties (e.g. transition strengths 2022A&A...661A.140M, 2021MNRAS.508.2236B), have led to a significant revision of solar abundances. Initial results based on these more sophisticated methods favored a markedly lower solar metallicity [2009ARA&A..47..481A], particularly in the CNO elements, than obtained in the 90s from older techniques [1993oee..conf...15G, 1998SSRv...85..161G].
@@
The degeneracy between solar composition and opacities can be broken using CNO solar neutrinos, e.g., following the methodology developed in 2008ApJ...687..678H. Such a study was recently carried out by the Borexino collaboration [2020Natur.587..577B, 2022PhRvL.129y2701A, 2023PhRvD.108j2005B]. The Borexino measurement of the CNO-I cycle neutrino flux (the 13N and 15O fluxes) were used to determine the C+N core abundance. Results show a tension with LZ metallicity determinations, while being in better agreement with HZ mixtures. While the error budget is presently dominated by the uncertainty of the Borexino CNO neutrino measurement, a significant contributor to the error () is nuclear, due to uncertainties in , , and . The interpretation of future improved CNO neutrino measurements will be impacted, unless these nuclear physics uncertainties are reduced.
@@ II.1.2 The Gallium anomaly
The precisely known rate of electron capture on 71Ge to 71Ga, for which the half-life is 11.43(3) d [1985PhRvC..31..666H], places an important constraint on the neutrino absorption rate [2023PhLB..84237983G, PhysRevC.108.035502]. The half-life has now been remeasured by two groups, 2024PhRvC.109e5501N and PhysRevC.108.L021602, yielding values of 11.468(8) d and 11.46(4) d, consistent with the previous determination while increasing the precision. The allowed matrix element that governs neutrino capture to the ground state of 71Ge can be extracted to a precision of 1% from the electron capture rate. The ground-state transition by itself generates a significant anomaly: inclusion of the excited-state contributions doubles the effect. In Appendix A we describe some of the details of both the 1997 extraction of the allowed (Gamow Teller) strength by 1997PhRvC..56.3391B and the recent update of PhysRevC.108.035502, ELLIOTT2024104082.
@@ II.1.3 The Boron-8 neutrino spectrum
In 2011RvMP...83..195A it was decided to recommend the spectrum calculated by Winter et al. [2003PhRvL..91y2501W, 2006PhRvC..73b5503W, Win:2007] based on their measurement of the alpha spectrum from . This spectrum showed excellent agreement with an independent experiment [2006PhRvC..73e5802B]. Both radiative and recoil-order corrections were included by Winter et al. The radiative corrections are relatively small, due to a cancellation between the real and virtual contributions [1995PhRvD..52.5362B]. The recoil-order corrections are dominated by the weak magnetism part, which has been deduced from measurements of the analog electromagnetic decays [1995PhRvC..51.2778D].
Since then three additional measurements [2011PhRvC..83f5802K, 2012PhRvL.108p2502R, 2023PhRvC.107c2801L] of the -decay alpha spectrum have been performed. All three find that the peak of the alpha spectrum appears about 20 keV lower than determined by Winter et al. Overall, these new measurements yield differences in the neutrino spectrum of 5% below MeV, the energy above which the contribution becomes small compared to the hep component. Figure 4 shows the 8B spectral uncertainties on the scale of the hep spectrum. With resolution effects considered, it will be difficult to extract the hep flux from solar neutrino measurements without further reduction in 8B uncertainties.
@@ II.2 Experimental program on solar neutrinos
A broad range of technology can be used to interrogate solar neutrinos. Radiochemical experiments utilizing chlorine and gallium played a crucial role in the early days of the solar neutrino problem. Large monolithic detectors have achieved great success in real-time measurements – from the water Cherenkov detectors such as Kamiokande, SNO, and SuperK, which could use the unique topology of Cherenkov light to point back to the neutrinos’ origin, to liquid scintillator detectors such as KamLAND and Borexino, whose high light yield and low threshold allow for precision spectroscopy.
@@ II.2.1 Super-Kamiokande and Hyper-Kamiokande
The first real-time detection of solar neutrinos was achieved by the Kamiokande experiment [1989PhRvL..63...16H], which detected the neutrinos via elastic scattering of electrons, a process in which the outgoing electron’s direction is highly correlated to that of the incoming neutrino. This allowed Kamiokande to directly point back to the neutrinos’ origin: our Sun. The successor experiment, Super-Kamiokande, is a 50-kton volume of pure water, surrounded by over 10,000 photon detectors, which has achieved an unparalleled program of neutrino observations and other physics over its several decades of operation. This program has included the highest-precision measurement of the elastic scattering signal from solar neutrino interactions, as well as sensitive searches for the day/night distortions of the spectral shape [2016PhRvD..94e2010A]. Now filled with a gadolinium additive to enhance neutron capture, SuperK’s primary current focus is the search for the Diffuse Supernova Neutrino Background (DSNB) via inverse beta decay [2023ApJ...951L..27H]. It also continues its atmospheric neutrino program: SuperK measurements made 25 years ago demonstrated the neutrino oscillations were responsible for the puzzling zenith-angle-dependence of this flux [PhysRevLett.81.1562].
@@ II.2.2 Sudbury Neutrino Observatory and SNO+
SNO’s unique use of heavy water as a target medium offered the ability to detect solar neutrinos via two additional interactions besides elastic scattering. In the charged current interaction an electron neutrino interacts with the neutron in the deuteron, producing an electron and a proton. This interaction path is sensitive only to electron-flavor neutrinos at the few-MeV energy scale of solar neutrinos, providing a measurement of the pure flux. SNO also had access to the neutral current interaction, in which a neutrino of any flavor interacts with the deuteron, breaking it apart into its constituent nucleons. Being equally sensitive to all active flavors, the neutral current interaction yields a measurement of the total, flavor-blind neutrino flux. It was this capability that allowed SNO to resolve the solar neutrino problem, demonstrating that most of the s produced by the Sun were arriving at earth in a different flavor state [2001PhRvL..87g1301A, 2002PhRvL..89a1301A, 2004PhRvL..92r1301A].
@@
SNO ceased data taking in 2006. Analysis of the data set has continued, and produced a number of new results, including constraints on non-standard effects such as Lorentz violation and neutrino decay [2018PhRvD..98k2013A, 2019PhRvD..99c2013A], limits on the fluxes of hep neutrinos and the diffuse supernova neutrino background (DSNB) [2020PhRvD.102f2006A], and neutron production from cosmogenic muons and from atmospheric neutrino interactions [2019PhRvD..99k2007A, 2019PhRvD.100k2005A].
@@
After decommissioning, the detector was repurposed for the SNO+ experiment, in which the target material was replaced with a pure organic liquid scintillator: linear alkyl benzene (LAB), loaded with 2.2 g/L of the secondary fluor, PPO. The high light yield of this scintillator, coupled with the location in SNOLAB –at 6-km water equivalent, one of the deepest underground laboratories in the world– and an extremely well-understood detector, allow for a range of high-precision measurement programs. In a preliminary water phase, SNO+ demonstrated detection of neutron captures on hydrogen [PhysRevC.102.014002], an impressive technical achievement in an unloaded water detector; a low background measurement of 8B solar neutrinos [PhysRevD.99.012012]; several searches for invisible modes of nucleon decay [PhysRevD.99.032008, PhysRevD.105.112012]; and detection of antineutrinos from reactors over 200 km away [PhysRevLett.130.091801]. The low backgrounds achieved and the detector’s demonstrated technical capabilities set the stage for future measurements. In a “partial-fill” stage, in which the upper half of the detector was filled with liquid scintillator while the lower half still contained water, the first demonstration of event-level direction reconstruction was achieved for 8B neutrinos in a scintillator detector [PhysRevD.109.072002]. This was facilitated by a lower loading of PPO at that time, which results in lower scintillation yield and a slower time profile, effectively enhancing the Cherenkov component, which can be leveraged for directional information.
@@
Now fully filled with liquid scintillator, SNO+ is the deepest, largest operating liquid scintillator detector in the world. The future program will include measurements of several solar neutrino fluxes, as well as the 8B neutrino energy spectrum.
@@ II.2.3 Borexino
Water Cherenkov experiments such as Kamiokande, SuperK and SNO provided the first real-time measurement of solar neutrinos but their relatively high energy threshold made them sensitive only to a small fraction of the solar neutrino flux. To study in real time the bulk of solar neutrino emission a different detector technology was required. The Borexino experiment, located deep underground at the INFN Laboratori Nazionali del Gran Sasso, used an organic liquid scintillator target made of pseudocumene with PPO to detect the elastic scattering of solar neutrinos off electrons. The high light yield of the liquid scintillator made it possible to significantly lower the energy threshold, but given the small signal rate and the lack of a clear signature to separate it from the background (such as the direction indicated by the emission of Cherenkov radiation), the experiment required a long preparatory phase to develop the most advanced techniques to suppress the background, especially the one due to radioactive contamination of the liquid scintillator itself.
@@
The detector’s extreme radiopurity was key to its success, and over its 14 years of data taking, ending in October 2021, Borexino proved itself capable of covering the entire solar neutrino spectrum. At the beginning of its data taking in 2007 the level of 238U and 232Th contamination in Borexino were lower than and respectively [2008PhLB..658..101B], paving the way for the first measurement of the 7Be sub-MeV solar neutrinos, followed by a low-threshold measurement of the 8B flux and a first indication of the pep neutrinos. After a purification campaign that further reduced the liquid scintillator contamination, Borexino aimed at improving the accuracy of its results in the measurement of the pp-chain solar neutrinos. The flux of neutrinos produced in the pp fusion process was first reported in 2014Natur.512..383B and later improved, along with all the other fluxes from the pp chain, apart from the hep neutrinos [2018Natur.562..505B, 2019PhRvD.100h2004A, 2020PhRvD.101f2001A]. The pp flux was determined to an uncertainty of 11%, improving the neutrino-based estimate of solar luminosity, while 7Be neutrinos were measured with an uncertainty of , half that of the SSM prediction. Furthermore, using the SSM to constrain the flux of CNO neutrinos, the pep neutrino signal was established with a significance exceeding for the first time, and the 8B flux was measured with a threshold as low as 3 MeV.
@@
After completing the investigation of the pp-chain neutrinos, Borexino reported the first detection of solar neutrinos produced in the CNO-cycle [2020Natur.587..577B], demonstrating directly that this mode of hydrogen burning operates in stars. Profiting from a larger exposure and a better understanding of the radioactive backgrounds enabled by the unprecedented thermal stability of the detector, Borexino further improved its measurement of the CNO neutrino flux [2022PhRvL.129y2701A, 2023PhRvD.108j2005B], where the new CNO result was used in combination with existing solar neutrino results to probe solar composition.